OpenStudy (anonymous):

there are 6 chemical elements called NOBLE GASES.suppose you write the names of all 112 elements on cards and select a card without looking .what is the probability of not picking a NOBLE GAS

OpenStudy (anonymous):

Probability of occurence=P(E) P(E)=6/112=3/56 Probability of non occurence=1-P(E) =1-3/56 =(56-3)/56 =53/56=0.95